if sinx+sin 2 x=2 then find,cos 2 x+cos 4 x

if sinx+sin2x=2


then find,cos2x+cos4x

Grade:10

2 Answers

Vinay Arya
13 years ago

Hi Vishwa Sharma

sinx+sin2x=2

=>sinx+sin2x-2=0

=>(sinx-1)(sinx+2)=0

=>sinx=1 or sinx=-2

 We know that sinx cannot be equal to -2.Hence rejaecting sinx=-2

 Therefore, sinx=1 =>x=pi/2

  cos2x+cos4x=xos2pi/2 + cos4pi/2 =0

 

Chaitanya Borah
13 years ago

sinx+sin2x=2

=>sinx=1+1-sin2x

=>sinx=1+cos2x

=>cos2x=sinx-1

now,

cos2x+cos4x

=cos2x+(sinx-1)2

=cos2x+sin2x-2sinx+1

=1-2sinx+1

=2(1-sinx)